# Destruction of ’Peas in a Pod?’ A Candidate Multi-planet System Around the Nearby, Bright Star, HD208487

### v1

@@ 6.1 Fitting Stellar Activity with EMPEROR
Table 6: EMPEROR results for Combined FWHM for HD208487 with $1\sigma$ errors.
Activity Index: Combined FWHM (ecc 0.3 prior)
k1 k2
@@ 7.1 ASAS Photometry
Figure 7: Left (Right): ASAS V-band photometry (Hipparcos Photometry). Top: ”A” and ”B” flagged V-band ASAS photometry (all Hipparcos data) for HD208487. Middle: Cleaned V-band ASAS photometry (Hipparcos) for HD208487. Bottom Left: Generalized Lomb-Scargle periodogram of the photometry. Horizontal lines correspond to the 0.1% (dotted blue) FAP significance level computed via 5000 bootstrap iterations on the photometry time series. The vertical green lines represent the periods of the three planets.
@@
7.1.1 Fitting Stellar Cycles
@@ Acknowledgements.
Appendix A Complete Radial Velocity and Stellar Activity Index Data
The following tables list all of the RVs and stellar activity index data utilized in this paper.
Table 8: Complete AAT data
@@ Appendix A Complete Radial Velocity and Stellar Activity Index Data
2451828.93670 11.580 4.030
2452062.28449 -9.140 2.350
2452092.18604 4.530 2.540
2452130.09003 20.530 2.830
2452151.06087 16.690 2.690
2452151.99385 12.150 2.330
2452152.97460 16.090 2.160
2452154.05537 3.020 3.310
2452186.93709 -16.560 2.120
2452188.01333 -26.390 2.180
2452189.07521 -17.590 1.980
2452189.95958 -15.190 2.420
2452388.30750 4.590 2.450
2452477.16326 1.440 3.070
2452598.94574 -0.560 2.690
2452784.26723 24.630 1.840
2452859.14214 2.890 2.470
2452943.06869 5.370 2.540
2452946.99543 -3.200 2.790
2453214.19050 -14.330 2.250
2453215.18430 -10.380 1.900
2453216.22254 -14.000 1.730
2453243.24872 -20.500 2.940
2453244.14557 -9.770 2.240
2453245.14036 -22.760 2.640
2453245.93542 -9.930 2.700
2453282.01051 3.900 5.840
2453489.27774 -12.640 1.950
2453573.18022 37.850 2.800
2453664.98472 13.860 1.950
2453945.22713 19.400 1.360
2454012.98659 -22.590 1.490
2454038.99785 -8.610 3.210
2454254.21331 -22.420 2.130
2454338.22561 21.420 2.370
2454371.09020 -1.740 1.700
2455103.01798 -0.130 2.050
2455399.21234 12.200 2.110
2455430.18814 -0.880 2.560
2455845.97594 12.180 2.830
2456467.27664 -14.020 2.400
2456935.06220 11.780 2.990
2456969.95622 -18.000 3.300
Table 9: Complete TERRA data
@@
2453864.88194 -34.095 0.691 0.164 8.465 19.668
2453865.92725 -30.038 0.553 0.164 8.457 20.273
2453866.90195 -31.316 0.777 0.164 8.464 19.119
2453867.92542 -35.551 0.762 0.162 8.464 22.462
2453868.91623 -34.276 0.702 0.162 8.463 18.725
2453869.88794 -36.017 0.893 0.161 8.463 18.353
2453870.87586 -35.859 0.427 0.159 8.463 17.593
2453882.92017 -35.249 0.824 0.160 8.465 20.386
2453883.90239 -33.331 0.622 0.160 8.460 17.748
2453886.91959 -34.368 0.611 0.161 8.466 21.496
2453887.90246 -28.890 0.698 0.161 8.467 16.819
2454254.91746 -37.187 1.115 0.161 8.469 18.714
2454255.82308 -37.086 0.891 0.163 8.463 19.590
2454257.90216 -30.784 1.004 0.165 8.462 22.762
2454258.93746 -31.501 1.006 0.167 8.472 18.799
Table 10: Complete PFS data
@@
2456876.82515 -23.179 1.897 0.144
2457199.84220 13.107 1.641 0.142
2457206.84952 3.411 1.813 0.140
2457259.72419 -7.563 2.018 0.153
2457321.58081 23.976 1.769 0.144
2457327.62340 22.557 1.601 0.144
2457529.89692 -10.381 1.543 0.143
2457554.89578 -1.735 1.763 0.146
2457616.64569 -29.852 1.633 0.140
2457620.64176 -24.033 2.066 0.339
2458266.84871 -14.930 1.449 0.150
2458272.91464 -11.629 1.389 0.145
2458293.80551 -7.507 1.776 0.148
2458297.82056 -2.482 1.591 0.145
2458298.91657 0.899 1.614 0.144
2458355.65345 26.318 1.910 0.157
2458355.65588 21.484 1.864 0.158
2458355.65830 27.853 1.945 0.157
2458625.87440 4.666 1.677 0.163
2458625.87721 6.568 1.821 0.162
2458625.88005 3.032 1.769 0.162
2458674.82066 -16.397 1.823 0.152
2458674.82665 -13.872 1.755 0.152
2458685.76083 -7.206 1.650 0.152
2458685.76531 -7.692 1.759 0.157
2458738.60957 16.696 1.888 0.159
2458738.61344 17.112 1.899 0.158
2459152.57287 -3.915 1.665 0.150
2459152.57582 -5.512 1.626 0.153
2459152.57882 -0.662 1.668 0.153
2459359.89486 3.947 1.529 0.158
2459359.89679 0.659 1.597 0.158
2459359.89874 4.639 1.687 0.156
2459359.90067 3.909 1.690 0.157
2459449.74696 -16.079 1.871 0.160
2459449.75193 -15.890 1.990 0.163
2459475.63281 -10.275 1.766 0.157
2459475.63674 -13.527 1.810 0.157
2459508.64032 1.264 1.723 0.161
2459508.64415 2.893 1.742 0.160
Table 11: Complete TERRA2 data
@@ Appendix B Split RV GLS Periodograms
Figure 10: Top Panel: Original GLS periodogram of the combined and mean-subtracted Low S-index RV data. Second Panel: GLS periodogram after fitting and removing the first signal. Third Panel: GLS periodogram after fitting and removing both the first and second signals. Bottom Panel: GLS periodogram after manually fitting the third signal and removing all three previously detected signals. The blue, dotted horizontal lines represents the 0.1% False Alarm Probability (FAP) level estimated from 5000 bootsrap resamplings.
Figure 11: Top Panel: Original GLS periodogram of the combined and mean-subtracted High S-index RV data. Second Panel: GLS periodogram after fitting and removing the first signal. The blue, dotted horizontal lines represents the 0.1% False Alarm Probability (FAP) level estimated from 5000 bootsrap resamplings.

### v3

@@ 6.2 Fitting stellar activity with EMPEROR
Table 6: EMPEROR results for combined FWHM with $1\sigma$ errors.
$P$ (days) K(m/s)K\ ($\mathrm${m/s}) $e$ BIC
@@ 7.1 ASAS photometry
Figure 7: Left (right): ASAS V-band photometry (Hipparcos photometry). Top: ”A” and ”B” flagged V-band ASAS photometry (all Hipparcos data) for HD208487. Middle: cleaned V-band ASAS (Hipparcos) photometry for HD208487. Bottom: GLS periodogram of the photometry. Horizontal lines correspond to the 0.1% (dotted blue) FAP significance level computed via 5000 bootstrap iterations on the photometry time series. The vertical green lines represent the periods of the three planets.
@@
7.2 Fitting ASAS stellar cycles
@@ 9 Conclusions
Data Availability
Tables A.1-A.4 are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
Appendix A RV and stellar activity index data
The following tables list the first five RVs and stellar activity index data for each instrument utilized in this paper. The full data can be found at the CDS.
Table 8: First five AAT data
@@ Appendix A RV and stellar activity index data
Table 9: First five TERRA data
@@
Table 10: First five PFS data
@@
Table 11: First five TERRA2 data
@@ Appendix B Split RV GLS periodograms
Figure 10: Top panel: original GLS periodogram of the combined and mean-subtracted low S-index RV data. Second panel: GLS periodogram after fitting and removing the first signal. Third panel: GLS periodogram after fitting and removing both the first and second signals. Bottom panel: GLS periodogram after manually fitting the third signal and removing all three previously detected signals. The blue, dotted horizontal lines represent the 0.1% FAP level estimated from 5000 bootstrap resamplings.
Figure 11: Top panel: original GLS periodogram of the combined and mean-subtracted high S-index RV data. Second panel: GLS periodogram after fitting and removing the first signal. The blue, dotted horizontal lines represent the 0.1% FAP level estimated from 5000 bootstrap resamplings.